A rectangle measures 4/7 inches by 3/5 inches. What is the perimeter of the book? (Use P= 2L + 2W) can you please assist me with this one?
Answer by rfer(16322)

P=2(4/7)+2(3/5)
P=2(20/35)+2(21/35)
P=40/35+42/35
P=82/35
P=2 12/35 in